OpenStudy (anonymous):

Which is not an equation of the line that passes through the points (1, 1) and (5, 5)?

OpenStudy (anonymous):

the equation of the line is y-y1=m(x-x1) 1) finding m=slope m=(y-y1)/(x-x1) 2) using the equation above y-y1=m(x-x1) choosing any points and sub this point into the equation x1 and y1 3) solve for y